Miriam is 10 years older than Ricardo. Two years ago, Miriam was 3 times as old as Ricardo is now. How old are they now?
Answer by jorel1380(3719)

R+10=M
3R=M-2
3(M-10)=M-2
3M-30=M-2
2M=28
M=14
R=4
Ricardo is 4. Miriam is 14..